OpenStudy (anonymous):

One person can do a certain job in ten minutes, and another person can do the same job in fifteen minutes. How many minutes will they take to do the job together? If x represents how many minutes to do the job together, then how much of the job does the slowest person do?

OpenStudy (anonymous):

1/10 + 1/15 = 1/y where y is time to do job together

hero (hero):

That still doesn't answer the question

hero (hero):

She already knows y = 6

OpenStudy (anonymous):

I don't get that... sorry!

OpenStudy (anonymous):

I still need help.

OpenStudy (anonymous):

well - its the way to do first part

OpenStudy (anonymous):

y = 6 is the solution of that equation

OpenStudy (anonymous):

These are my options: a) x/15 b) x/10 c) 15/x

hero (hero):

You could also have done it like this: \(\frac{M \times W}{M + W} = x\) \(\frac{10 \times 15}{10 + 15} = x\) \(10 \times 15 = 10x + 15x\) \(150 = 10x + 15x\) \(1 = \frac{10x + 15x}{150}\) \(1 = \frac{10x}{150} + \frac{15x}{150}\) \(1 = \frac{x}{15} + \frac{x}{10}\)

hero (hero):

so if we are going by the first person, which we know did the job in 10 minutes, then they must have only done x/15 of the job

hero (hero):

Remember this is fractions so \(\frac{x}{15} < \frac{x}{10}\)

hero (hero):

So it makes sense

OpenStudy (anonymous):

Okay...wow! You're good!

hero (hero):

Also, to check, since we know that x = 6, we can make sure that \(\frac{6}{15} + \frac{6}{10} = 1\)